OpenStudy (johnnewman):

What are the square roots of 196? A. 98 and –98 B. 14 and –14 C. 13 and –13 D. 16 and –16

pooja195 (pooja195):

Well what is the square root of 196 do you know?

OpenStudy (alexandervonhumboldt2):

\[\sqrt{196}=\sqrt{14*14}=?\]

OpenStudy (johnnewman):

ok

OpenStudy (alexandervonhumboldt2):

hint: 14*14=196 (-14)*(-14)=196 as well

OpenStudy (johnnewman):

ok thanks

OpenStudy (johnnewman):

thanks pooja195

pooja195 (pooja195):

yw :) Welcome to openstudy!

OpenStudy (johnnewman):

I am new thanks

pooja195 (pooja195):

^-^ hmm \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (johnnewman):

it my frist time here

pooja195 (pooja195):

I see I hope you have a great time here :)